OpenStudy (anonymous):

what is the LCM of 3,9,15?

OpenStudy (geometry_hater):

what number will go into 3,9 and 15?

OpenStudy (anonymous):

its three!

OpenStudy (amistre64):

..... that comes across as a bit rude :/

OpenStudy (anonymous):

what is rude

OpenStudy (anonymous):

3 is the answer :)

OpenStudy (amistre64):

i was sitting there, taking my time to answer your questions, and you simply walk away and start up another one ... no thank you or nothing ....

OpenStudy (amistre64):

LCM aint 3, the GCD is 3 ....

OpenStudy (anonymous):

then 1..?

OpenStudy (geometry_hater):

No its not 1

OpenStudy (amistre64):

factor each term into its primes, then see what you need to "add" to each one to make a single value

OpenStudy (anonymous):

well then i dont know.

OpenStudy (amistre64):

3 = 3 9 = 3*3 15 = 3*5 from this we can see that: the 3 needs a 3 and 5 the 9 needs a 5 and the 15 needs a 3 that way they are all: 3*3*5

OpenStudy (anonymous):

most of the time 1 is not accepted so i would say the next one is 3

OpenStudy (anonymous):

yeah thats what i thought too! its three cause that is the smalest # that goes into all thoes other #s

OpenStudy (amistre64):

... how about we do some actual math instead of making up stuff :/

OpenStudy (amistre64):

it aint 3, 9 * ? = 3 15 * ? = 3